OpenStudy (anonymous):

I need the solutions of 8= (x+3)^2

OpenStudy (anonymous):

Lets first F.O.I.L. the (x+3)^2

OpenStudy (anonymous):

We get \[x ^{2} + 6x + 9\]

OpenStudy (anonymous):

So we have \[8 = x ^{2}+6x+9\]

OpenStudy (anonymous):

Let's subtract the 9 from both sides.

OpenStudy (anonymous):

-1 = x^2 + 6x

OpenStudy (anonymous):

i need help please help me

OpenStudy (anonymous):

Sorry @verity, ask your question elsewhere. I am busy here.

OpenStudy (anonymous):

No offense!

OpenStudy (anonymous):

@nodive Do you have any questions so far?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Great!!

OpenStudy (campbell_st):

take the square root of both sides \[\pm \sqrt{8} = x + 3\] now subtract 3 \[-3 \pm \sqrt{8} = x\] the radical can be simplified if necessary.... hope it helps

OpenStudy (anonymous):

So from this point. We want to solve for x. -1 = x^2 + 6x Let's factor out the x. -1 = x (x+6)

OpenStudy (anonymous):

Sorry. Step back a few steps.

OpenStudy (campbell_st):

if you distribute and then set the equation to zero, you will need the general quadratic formula to get the solution. by starting with doing the opposites it avoids unnecessary work.

OpenStudy (anonymous):

Omg, I am really sorry, @campbell_st is absolutely right. I apologize so much for walking you through wrong..

OpenStudy (anonymous):

g(x)=x^2+6x+1 0=x^2+6x+1 -1=x^2+6x -1+9=x^2+6x+9 8=x^2+6x+9 8= (x+3)^2

OpenStudy (campbell_st):

the method @Cosmichaotic has posted is correct... just the long way around

OpenStudy (anonymous):

\[x = -b \pm \sqrt{\frac{ b ^{2-4ac} }{ 2a }}\]

OpenStudy (anonymous):

This is the quadratic formula. Once you put the equation in ax^2 + bx + c = 0 (Standard Form), you will know where to substitute... I can walk you through this if you like.

OpenStudy (anonymous):

8 = (x+3)^2 -> (x+3)^2 = 8 -> x^2 + 6x + 9 = 8 -> x^2 + 6x + 9 - 8 = 0 -> x^2 + 6x + 1 = 0 Where, a = 1 b = 6 c = 1

OpenStudy (anonymous):

Lol, I am really destroying this answer aren't I....

OpenStudy (anonymous):

so \[-6\pm 4\]

OpenStudy (campbell_st):

nope if you use the general quadratic formula its \[\frac{-6 \pm \sqrt{6^2 - 4 \times 1 \times 1}}{2 \times 1} = \frac{-6 \pm \sqrt{32}}{2}\] now simplify

OpenStudy (anonymous):

Well done. Sorry about that.

OpenStudy (anonymous):

i couldnt have done any better @Cosmichaotic

OpenStudy (campbell_st):

to simplify it becomes \[\frac{-6 \pm \sqrt{4 \times 8}}{2} = \frac{-6 \pm \sqrt{4}\times \sqrt{8}}{2} = \frac{-6 \pm 2\sqrt{8}}{2}\] then you'll get \[-3 \pm \sqrt{8}\] as I said... the long way round

OpenStudy (anonymous):

thank yall!

OpenStudy (tkhunny):

One of the solutions presented made no sense. It was this one: \(8 = (x+3)^{2}\) Take the Square Root of both sides. \(\pm \sqrt{8} = x+3\) This is no good. \(\sqrt{8} > 0 \) What should have happened was: Take the Square Root of both sides. \(\sqrt{8} = \pm (x+3)\) It may not make much difference as far as the solution is concerned, but it presents a better understanding of the square root operation.

OpenStudy (anonymous):

tbh im just lost in general

OpenStudy (tkhunny):

There are multiple examples. Please read carefully. You should see it.